If μ = 9 ounces, σ = 0.5 ounces, and n = 9, calculate the 3-sigma control limits for the x-bar chart.
1
Expert's answer
2020-12-06T20:09:44-0500

For 3-sigma control limits for the x-bar chart Z = 3.

"UCL = μ + Z\frac{σ}{\sqrt{n}} \\

= 9 + 3\frac{0.5}{\sqrt{9}} \\

= 9.5 \\

LCL = μ - Z\frac{σ}{\sqrt{n}} \\

= 9 - 3\frac{0.5}{\sqrt{9}} \\

= 8.5"

Answer: 8.5 to 9.5 ounces